OpenStudy (anonymous):

How many solutions does this system have? 2x+y=3 6x=9-3y one none infinite two

hero (hero):

2x+y=3 6x=9-3y y = 3-2x y = 3-2x infinite solutions

hero (hero):

They're both the same graph

OpenStudy (anonymous):

 haha :P the graph should look something like this